# Stretching Beyond the Obvious: A Gradient-Free Framework to Unveil the Hidden Landscape of Visual Invariance

### v1

@@ B.1 Statistical Analysis for Classification Experiments
Analysis of data from classification experiments focused on identifying accuracy differences between 9 image types: natural images, MEI’s from robust and vanilla networks, and invariant images from 3 different layers for both network types. We fitted a generalized linear mixed model (GLMM) with a logistic link function for binary correct vs. incorrect trial responses, including random intercepts for participants and semantic image categories. In addition to data from human participants (Fig.5b, top), we applied the same methodology to separately analyze AlexNet’s responses to the full complement of 1080 images (see Section 3.4), but with GLMM random intercepts for semantic image categories only (Fig.5b, bottom).
@@
For both human participants and AlexNet, We used an omnibus test (type II Wald chi-square) to assess overall differences among the 9 image types, followed by 8 planned contrasts via estimated marginal means: (i) robust vs. vanilla MEIs, (ii) robust vs. vanilla invariant images averaged across the 3 representation layers, and (iii-viii) within each network type, all 3 possible pairwise comparisons between invariant types at different layers. Our analysis employed the Benjamini-Hochberg procedure to control the false discovery rate under multiple comparisons [53]. The results of the planned comparisons for both humans and AlexNet are provided in Table 1. Notably, the analysis of human data included a larger number of data points overall: while each human only viewed 540 images as opposed to the AlexNet’s 1080, there were 25 human participants responding independently unlike the single response per image from AlexNet.
Table 1: Statistical results from omnibus test (type II Wald chi-square) and post-hoc pairwise comparisons to test differences among stimulus types in the human/model image classification experiments. All $p$ values are adjusted using the Benjamini-Hochberg correction (separately for humans and AlexNet), with significant values in bold.
Human subjects AlexNet
Comparison z-value p-value z-value p-value
Omnibus test (effect of stimulus type) - <0.0001 - <0.0001
Average Robust Stretch vs Average Standard Stretch 36.63 <0.0001 10.02 <0.0001
Robust layer4_conv7 vs Robust layer3_conv1 -4.94 <0.0001 1.06 0.39
Standard layer4_conv7 vs Standard layer3_conv1 -8.41 <0.0001 7.28 <0.0001
Robust layer4_conv7 vs Robust pixel space -9.52 <0.0001 -0.63 0.60
Standard layer4_conv7 vs Standard pixel space 7.49 <0.0001 5.44 <0.0001
Robust layer3_conv1 vs Robust pixel space -4.66 <0.0001 -1.68 0.15
Standard layer3_conv1 vs Standard pixel space -0.99 0.32 -2.43 0.030
Robust MEI vs Standard MEI 32.96 <0.0001 0.22 0.83
Appendix C SnS is effective in targeting units in hidden layers
@@
Figure S3: SnS is effective with targets in hidden layers (a) Comparison of MEIs (top row) with SnS-generated invariant images (bottom row) for a representative target unit in Layer3_conv1 of robust ResNet50. SnS images, synthesized to be invariant in pixel space, were generated using the green-highlighted MEI as a reference. Red rectangles identify the unit’s key responsive regions. This comparison highlights SnS’s ability to distill the core visual features a unit detects, offering clearer insight than MEIs alone. (b-d) The same SnS-based feature visualization approach (as in (a) is applied to different representative target units: (b) Layer3_conv1 of a standard ResNet50; (c) Layer4_conv7 of a robust ResNet50; and (d) Layer4_conv7 of a standard ResNet50. (e) Final activation distributions for the target units from the SnS optimization in hidden layers. As in Fig. S1, convergence performance is compared to natural image statistics
@@
To investigate this, we conducted SnS invariance experiments targeting 50 distinct units in both layer3_conv1 and layer4_conv7 of the ResNet50 network, using pixel space as the input representation. For each unit, we performed 10 optimization runs initiated with different random seeds. In the standard network, SnS terminated by early stopping in 67.20% of runs at layer3_conv1 and 27.40% at layer4_conv7; in the robust network, convergence was 87.80% and 89.40%, respectively. Importantly, even in non-converging instances, the optimization process often guided the input towards stimuli that elicited high unit activations, consistent with functionally relevant regimes identified through natural image statistics on ImageNet (Fig. S3e).
@@
Qualitatively, the SnS-generated invariant images served as effective “feature visualizers.” This was particularly insightful as the precise selectivity of these hidden units is unknown a priori. The resulting visualizations clearly delineated the specific image features to which a target neuron was most sensitive (region of interest (ROI)), effectively isolating these core features from irrelevant contextual details (Fig. S3a-d (bottom rows)). This level of feature disentanglement and clarity surpassed that observed in images optimized by methods like XDREAM [10], which often retain more complex, less isolated visual elements (Fig. S3a-d (top rows)). Moreover, by comparing the reference image to multiple invariant instances (i.e. initialized with multiple random seeds), the images revealed specific transformations the identified ROI could tolerate while preserving the unit’s activation level, offering insights into the unit’s invariances (Fig. S3a-d (bottom rows)).
Appendix D SnS is robust to subsampling of the representation space
@@
Figure S4: SnS invariance is robust to subsampling of the representation space (a) Example MEI and associated invariant images obtained for a readout neuron (”candle”) in robust ResNet-50 stretching the mid_level layer (i.e. Layer3_conv1). Each row represents examples for different levels of subsampling: using the full layer ($1^{st}$ row), subsampled spaces of size $n=1000$ and $n=100$ ($2^{nd}$ and $3^{rd}$ rows respectively). Multiple results are shown for different random initialization seeds (columns). (b) Normalized average distance between the invariant images generated by SnS (stretching in Layer3_conv1) and their reference MEIs across the different stages of a robust ResNet50. Different lines indicate different cardinalities of units in the representation space used for optimization (i.e. all, $n=1000$ and $n=100$). The average distance between multiple MEIs generated via XDREAM is reported for comparison. (c-d) Same as (a) and (b) respectively, but stretching was performed on robust ResNet-50 in the high_level layer (i.e. Layer4_conv7).
@@
SnS, as a gradientless and model-agnostic method, offers significant advantages for neuroscience research, particularly in scenarios where constructing detailed “digital twin” models (as described in Section 2) is impractical. However, a critical challenge for applying SnS to study hierarchical invariances in the brain is the inherent undersampling of representational spaces. Despite recent advances in electrophysiology and calcium imaging enabling simultaneous recording from hundreds or thousands of neurons [54, 55], these recordings invariably capture only a fraction of the total neural population within a given brain area.
@@
To evaluate SnS’s performance under conditions analogous to experimental neuroscience, we investigated how subsampling the representational space impacts the qualitative nature and quantitative characteristics (assessed via representational distance analysis) of the invariances identified by SnS. Specifically, we focused on the same 77 readout target units previously analyzed in Section 4.2. For these units, we identified invariances after subsampling their corresponding representational spaces (either layer3_conv1 or layer4_conv7 of the CNN) to either 1000 or 100 randomly selected units. To ensure robustness, this process was repeated 10 times for each target unit and subsampling condition, using different random selections of representational space units. As a reference, 1000 units represent 0.5% of the units in layer3_conv1 ($n=200704$), while they constitute 4% of layer4_conv7 ($n=25088$).
@@
A plausible explanation for this robustness to subsampling lies in the inherent representational redundancy present in both biological neural circuits and artificial neural networks. In biological systems, information is often encoded in a distributed manner across populations [56] [57], where multiple neurons may carry similar or overlapping information, contributing to robustness against noise and cell loss [58]. Similarly, deep neural networks, including CNNs, have been shown to learn redundant features or possess over-parameterization, where multiple units or weights contribute to similar computations [59]. Consequently, a sufficiently large and representative subsample of units can still capture the dominant computational characteristics and drive the overall representation into similar regimes as the full layer, explaining the observed consistency in identified invariances.

### v2

@@ B.1 Statistical Analysis for Classification Experiments
Analysis of data from classification experiments focused on identifying accuracy differences between 9 image types: natural images, MEI’s from robust and standard networks, and invariant images from 3 different layers for both network types. We fitted a generalized linear mixed model (GLMM) with a logistic link function for binary correct vs. incorrect trial responses, including random intercepts for participants and semantic image categories. We used an omnibus test (type II Wald chi-square) to assess overall differences among the 9 image types, followed by 8 planned contrasts via estimated marginal means: (i) robust vs. standard MEIs, (ii) robust vs. standard invariant images averaged across the 3 representation layers, and (iii-viii) within each network type, all 3 possible pairwise comparisons between invariant types at different layers. Our analysis employed the Benjamini-Hochberg procedure to control the false discovery rate under multiple comparisons (Benjamini and Hochberg, 1995). The results of the planned comparisons are provided in Table S2.
Figure S5: SnS generates effective adversarial and invariant examples for natural images (a) As in Fig. 2, the ordinate axis represents the activation reduction of a readout unit of $L_{2}$-robust ResNet50 with respect to the activation produced by a reference natural image, when SnS is applied to synthesize either adversarial or invariant images, while the abscissa represents the $L_{2}$ pixel distance between these synthetized images and the reference. In this experiment, differently from Fig. 2, the reference images were not the MEIs of the readout units but natural images belonging to the specific categories encoded by the units. Each data point reports the average ($\pm$ SD) of these metrics across $12\times 10$ image synthesis experiments, where $12$ are the targeted readout units and $10$ are the different reference images selected for each unit. (b) Examples of invariant and adversarial images obtained for 5 readout units of $L_{2}$-robust ResNet50, when the reference images were either their MEIs (top), as for the experiment shown in Fig. 2, or natural images (bottom), as for the experiment shown in (a).
Appendix C SnS generates effective adversarial and invariant examples for natural images
We replicated the experiment reported in Section 4.1 (see Fig. 2) also in the case where, for a target readout neuron, a highly effective natural image representative of the class encoded by the unit was used as reference for SnS optimization instead of a MEI. As for Section 4.1, both invariant and adversarial images were synthesized by stretching the pixel-level representation. In these experiment, we considered the $12$ readout units of robust ResNet50 encoding the $12$ image categories used for the interpretability experiment with human observers (see Sections 3.2, 4.3 and Supplementary Section B). For each unit, both the adversarial and invariant experiment was repeated for 10 different reference natural images, the same ones used for the interpretability experiment (for a total of $12\times 10$ invariant images and $12\times 10$ adversarial images). Example images are shown in Fig. S5b, bottom. Fig. S5a shows that, similarly to the experiment performed with MEI references (see Fig. 2), both adversarial and invariant images reached the desired functional activation regimes (i.e., for adversarial images: silencing the target unit without distancing too much from the reference image; for invariant images: departing substantially from the reference natural image, while at the same time preserving the same level of activation).
Appendix D Single-loss ablation experiments
To assess the functional contribution of each component of the SnS dual-loss objective (see Section 3.1 for a detailed description), we performed a set of ablation experiments in which we optimized images using only one loss term at a time. We repeated the experiment shown in Fig. 2, in which a readout unit is selected as the optimization target and pixel space is used as the representation space. For each target unit, we synthesized images under four single-loss variants: 1) Squeeze-up: minimize the target unit activation distance between the synthesized image and the reference MEI; 2) Stretch-up: maximize the activation distance from the reference MEI; 3) Squeeze-down: minimize the distance of the synthesized image from the reference MEI in the pixel representation; and 4) Stretch down: maximize such distance.
We used the same 77 readout units and 10 initialization seeds per unit as in the main experiments, resulting in 770 runs per single-loss condition. Squeeze-up and stretch-down followed the initialization and Pareto-front sorting used for SnS invariances; squeeze-down and stretch-up followed the procedures used for SnS adversarial optimizations (see Sections A.3 and 3.1).
Across all four variants, optimizing only one objective consistently failed to drive the other metric (activation distance or pixelspace distance) to its required regime (Fig. S6). None of the single-loss optimizations produced solutions that simultaneously exhibit the trade-off characteristic of SnS invariant or adversarial synthesis. This demonstrates that the dual-loss formulation is necessary for producing meaningful invariances and adversarial examples. Although the single-loss variants did not reproduce SnS behavior, each produced a characteristic synthesis regime that sheds light on the role of its loss component:
  • Squeeze-up: Optimizing only the squeeze term yielded images whose activation approached the invariant regime (i.e., close to the MEI’s activation), but whose pixel-space distance remained near the distance of the initial random seeds (see the black curve in Fig S2b, right when iterations is equal to 0). This behavior mirrors repeated XDREAM optimization with multiple initializations, producing multiple high-activation images that differ from each other but do not explore the far boundary of the invariance manifold. The explicit maximization of representation distance (e.g., in pixel space) is the factor that drives invariant images far away from its reference MEI, exploring the maximal invariant transformations within that representation space.
  • Stretch-up: Maximizing the activation distance naturally pushed images to diverge in pixel space as well. The optimization suppressed the target unit by making the image increasingly dissimilar from the reference. This behavior is essentially the opposite of SnS adversarial synthesis, where the unit is silenced with minimal pixel-space change.
  • Squeeze-down: Matching the reference MEI in pixel space also preserved the high activation of the reference. This optimization is notably similar to that used for model metamers (Feather et al., 2023), which enforce representation matching at internal layers. Thus, metamer synthesis emerges as a special case of the SnS loss structure when the squeeze term is applied alone to the representation space.
  • Stretch-down: Maximizing pixelspace distance alone caused images to diverge visually but produces no systematic change in target-unit activation. Indeed the activation of the target unit remained close to its initial value at the beginning of the optimization (see the black curve in Fig. S2b, left; iteration = 0), which is very far from the one produced by the MEI. This confirms that stretch-down alone does not find images relevant to the unit of interest.
Figure S6: Single-loss ablation experiments demonstrate the need of SnS dual loss. Comparison between SnS adversarial and invariance experiment distances (same as in Fig. 2) with the results of single loss optimizations (squeeze up, squeeze down, stretch up, stretch down). Each point corresponds to the mean $\pm$ SD of 770 experiments (i.e. 77 readout units x 10 random initializations per each unit). The black dotted line represents the median image $L_{2}$ distance among Imagenet images.
In summary, the single-loss ablations showed that none of the individual terms is sufficient to generate the invariant or adversarial images produced by SnS. The dual-loss structure is essential, with each component stabilizing and constraining the other. These results further validated SnS as a principled framework for probing model invariances and adversarial directions.
Appendix E SnS is effective in targeting units in hidden layers
@@
Figure S7: SnS is effective with targets in hidden layers (a) Comparison of MEIs (top row) with SnS-generated invariant images (bottom row) for a representative target unit in Layer3_conv1 of robust ResNet50. SnS images, synthesized to be invariant in pixel space, were generated using the green-highlighted MEI as a reference. Red rectangles identify the unit’s key responsive regions. This comparison highlights SnS’s ability to distill the core visual features a unit detects, offering clearer insight than MEIs alone. (b-d) The same SnS-based feature visualization approach (as in (a)) is applied to different representative target units: (b) Layer3_conv1 of a standard ResNet50; (c) Layer4_conv7 of a robust ResNet50; and (d) Layer4_conv7 of a standard ResNet50. (e) Final activation distributions for the target units from the SnS optimization in hidden layers. As in Fig. S2a, convergence performance is compared to natural image statistics
@@
To investigate this, we conducted SnS invariance experiments targeting 50 distinct units in both layer3_conv1 and layer4_conv7 of the ResNet50 network, using pixel space as the input representation. For each unit, we performed 10 optimization runs initiated with different random seeds. In the standard network, SnS terminated by early stopping in 67.20% of runs at layer3_conv1 and 27.40% at layer4_conv7; in the $L_{2}$-robust network, convergence was 87.80% and 89.40%, respectively. Importantly, even in non-converging instances, the optimization process often guided the input towards stimuli that elicited high unit activations, consistent with functionally relevant regimes identified through natural image statistics on ImageNet (Fig. S7e).
@@
Qualitatively, the SnS-generated invariant images served as effective “feature visualizers.” This was particularly insightful as the precise selectivity of these hidden units is unknown a priori. The resulting visualizations clearly delineated the specific image features to which a target neuron was most sensitive (region of interest (ROI)), effectively isolating these core features from irrelevant contextual details (Fig. S7a-d (bottom rows)). This level of feature disentanglement and clarity surpassed that observed in images optimized by methods like XDREAM (Xiao and Kreiman, 2020), which often retain more complex, less isolated visual elements (Fig. S7a-d (top rows)). Moreover, by comparing the reference image to multiple invariant instances (i.e. initialized with multiple random seeds), the images revealed specific transformations the identified ROI could tolerate while preserving the unit’s activation level, offering insights into the unit’s invariances (Fig. S7a-d (bottom rows)).
Appendix F SnS is robust to subsampling of the representation space
@@
Figure S8: SnS invariance is robust to subsampling of the representation space (a) Example MEI and associated invariant images obtained for a readout neuron (candle) in $L_{2}$ robust ResNet50 stretching the mid_level layer (i.e. Layer3_conv1). Each row represents examples for different levels of subsampling: using the full layer ($1^{st}$ row), subsampled spaces of size $n=1000$ and $n=100$ ($2^{nd}$ and $3^{rd}$ rows respectively). Multiple results are shown for different random initialization seeds (columns). (b) Normalized average distance between the invariant images generated by SnS (stretching in Layer3_conv1) and their reference MEIs across the different stages of a robust ResNet50. Different lines indicate different cardinalities of units in the representation space used for optimization (i.e. all, $n=1000$ and $n=100$). The average distance between multiple MEIs generated via XDREAM is reported for comparison. (c-d) Same as (a) and (b) respectively, but stretching was performed on robust ResNet50 in the high_level layer (i.e. Layer4_conv7).
@@
SnS, as a gradientless and model-agnostic method, offers significant advantages for neuroscience research, particularly in scenarios where constructing detailed “digital twin” models (as described in Section 2) is impractical. However, a critical challenge for applying SnS to study hierarchical invariances in the brain is the inherent undersampling of representational spaces. Despite recent advances in electrophysiology and calcium imaging enabling simultaneous recording from hundreds or thousands of neurons (Jun et al., 2017; Sofroniew et al., 2016), these recordings invariably capture only a fraction of the total neural population within a given brain area.
@@
To evaluate SnS’s performance under conditions analogous to experimental neuroscience, we investigated how subsampling the representational space impacts the qualitative nature and quantitative characteristics (assessed via representational distance analysis, see Section A.5.2 for methodological details) of the invariances identified by SnS. Specifically, we focused on the same 77 readout target units previously analyzed in Section 4.2. For these units, we identified invariances after subsampling their corresponding representational spaces (either layer3_conv1 or layer4_conv7 of the CNN) to either 1000 or 100 randomly selected units. To ensure robustness, this process was repeated 10 times for each target unit and subsampling condition, using different random selections of representational space units. As a reference, 1000 units represent 0.5% of the units in layer3_conv1 ($n=200704$), while they constitute 4% of layer4_conv7 ($n=25088$).
@@
A plausible explanation for this robustness to subsampling lies in the inherent representational redundancy present in both biological neural circuits and artificial neural networks. In biological systems, information is often encoded in a distributed manner across populations (Zohary et al., 1994) (Stringer et al., 2019), where multiple neurons may carry similar or overlapping information, contributing to robustness against noise and cell loss (Pouget et al., 2000). Similarly, deep neural networks, including CNNs, have been shown to learn redundant features or possess over-parameterization, where multiple units or weights contribute to similar computations (Denil et al., 2013). Consequently, a sufficiently large and representative subsample of units can still capture the dominant computational characteristics and drive the overall representation into similar regimes as the full layer, explaining the observed consistency in identified invariances.
Appendix G Hierarchical invariances in Vision Transformers
One of the strengths of SnS is that, being gradient-free, it can easily be applied (literally out-of-the-box) to virtually any visual processing architecture (biological or artificial), as long as one can measure the responses of its units to the images produced by SnS optimization algorithm. In cases where the goal is to uncover the invariance of the whole model to transformations of specific object classes, the additional constraint is to have access to the responses of the readout units for those classes. As such, SnS can directly be applied to investigate the invariance of readout units in vision transformers (ViTs).
To test this application, we ran the experiments of Section 3.2 (observer networks only) on a ViT-B-16 model pretrained on ImageNet and a CLIP-trained ViT-B-16 visual encoder (Radford et al., 2021). Since the CLIP model lacks explicit readout units, we trained a linear classifier atop the frozen visual encoder for ImageNet classification to construct class-specific detectors (top 1 validation accuracy: $78.94\%$). For both models, we considered the output linear layer of the MLP in the sixth encoder layer as mid-level layer and the one in the twelfth encoder layer as high level representation.
Our findings reveal that, similar to CNNs, the nature of the discovered invariances in ViTs depends on the hierarchical level of the stretched representation. Stretching the pixel space primarily altered low-level properties such as luminance, yielding invariances that were poorly interpretable (Fig. S9 a,b). Contrary to the experiments with ResNet50 (see Fig. 3 a,b), where a clear, qualitative distinction was observable between invariant images stretched in mid and late layers, in both ViTs we did not observe such difference (see Fig. S9 a,b). These images were also comparable in terms of invariance interpretability (Fig. S9 c). This observation is consistent with the view that ViTs learn less strictly hierarchical representations than CNNs, and that these models integrate information more globally (Dosovitskiy et al., 2020). Overall, there seems to be a little, but consistent across conditions advantage of the CLIP-trained version of ViT-B-16 in terms of interpretability. This trend seems to be similar between standard and robust network observers.
Figure S9: Invariances in vision transformers. (a) Example MEI and associated invariant images obtained for a readout neuron (cup) in standard ViT and computed for three different choices of the stretching representational stages (rows). Multiple results are shown for different random initialization seeds (columns) (b) Same as (a) but for a ViT encoder trained via CLIP. (c) Classification accuracies for ViT invariances across multiple robust (left) and standard (right) networks.
Appendix H SnS can be used to study invariance manifold geometry
SnS can be used to extensively sample the invariant manifold of a target unit, paving the way to the analyses of its intrinsic dimension (ID) and geometric structure. To show the feasibility of this approach we selected a single target readout neuron (the chickadee unit in the $L_{2}$ robust ResNet50 readout layer) and extensively sampled its maximally invariant images across three different stretching layers (the same ones used in Section 4.2): pixel space (input/low-level), Layer3_conv1 (mid-level), and Layer4_conv7 (high-level). For each stretching layer, we ran SnS 100 times for each of 5 different reference MEIs (obtained from independent XDREAM runs), yielding 500 invariant images per layer. To estimate the ID of each invariance manifold, we applied two distinct techniques to the vector representations of the invariant images: 1) PCA, measuring the number of components explaining 95% of the variance; and 2) two-nearest-neighbors intrinsic dimension estimation (2NN-ID; Facco et al. (2017)), a state-of-the-art, nonlinear estimator of intrinsic dimension that has been successfully applied to analyze image manifolds in both CNNs and visual cortical areas (Ansuini et al., 2019; Muratore et al., 2022a). 2NN-ID was computed for different subsamples of points (from $n=50$ to $n=500$ invariant images), repeating the process 50 times for each cardinality.
These two approaches returned dimensionalities in very different numerical ranges: PCA required hundreds of components to explain 95% of variance (from 230 to 340 PCs, Fig. S10a), whereas the ID estimated by 2NN-ID ranged between 10 and 30 across the three manifolds (Fig. S10b). This radical difference in ID estimation between a classic linear method (PCA) and 2NN-ID points to the fact that the structures of all 3 manifolds are highly nonlinear (Facco et al., 2017; Ansuini et al., 2019). This aligns with expectations for neurons high in the visual hierarchy whose responses abstract over complex real-world transformations.
Despite differences in scale, both methods agreed on a consistent ordering of dimensionality across stretching layers: lowest ID for pixel-level stretching, highest for mid-level stretching, and reduced ID again for high-level stretching. This pattern mirrors previously reported trends in the ID of image representations across deep CNNs (Ansuini et al., 2019), where ID initially increases, peaks in mid layers and then gradually diminishes in deeper layers. This result further reinforces the distinction between invariances uncovered at different stretching layers (Fig. 3), indicating the ability of SnS to discover axes of variation that are distinct not only in qualitative terms (Fig. 3a), but also in their dimensionality.
Figure S10: Hierarchical invariance manifolds differ consistently in their dimensionality. (a) PCA estimation of the dimensionality of the invariance manifolds obtained for an example $L_{2}$ robust ResNet50 readout unit by stretching the representations of the MEIs at different depths of the network: low-level (pixel space), mid-level (Layer3_conv1) and high-level (Layer4_conv7). The curves show the cumulative fraction of variance of each manifold (consisting of 500 sampled images) that was explained by considering an increasingly larger number of principal components (ranked according to their eigenvalue). For each stretching depth, the number of principal components necessary to explain 95% of data variance (horizontal black dashed line) is indicated by a color-specific vertical dashed line. (b) Dimensionality of the same invariance manifolds analyzed in a, but estimated using the nonlinear intrinsic dimension method 2NN-ID. The curves show the average intrinsic dimension $\pm$ SEM, as obtained for increasingly larger sets of invariant images sampled from the manifolds.
Appendix I SnS invariances and metamers are structurally and functionally distinct
To better contextualize SnS invariances within state-of-the-art techniques to probe invariances in neural networks, we carried out a systematic comparison between SnS-derived invariant images and gradient-based metamers. Metamers were generated using the original implementation of Feather et al. (2023), including the same hyperparameter settings mentioned in the original paper. Metamers were synthesized for the same $n=120$ natural images used in our behavioral experiment (Fig. 4), yielding for each natural image one SnS invariant and one metamer at each of three representational depths (pixel space, mid-level, high-level) in both standard and $L_{2}$-robust ResNet50 (for details see Section 3.1). First we compared metamers and SnS invariances in terms of interpretability. To do this, we replicated the experiment described in Fig. 4c with metamer stimuli. Using the same observer networks as in our main experiments (six architectures, each in both standard and $L_{2}$-robust versions; $n=12$ total (Section 4.3)), we recovered results consistent with Feather et al. (2023) (Table S3). Specifically, metamers produced by the $L_{2}$-robust ResNet50 remained highly interpretable across observer models at all synthesis depths (pixel space, layer 3, layer 4). In contrast, standard-network metamers showed the expected sharp decline in interpretability for late-layer metamers (layer 4) across all observer networks. Critically, this pattern is the opposite of what we observed for SnS hierarchical invariances (Fig. 4c; numerical values in Table S3). Indeed, for standard networks, late-layer SnS invariances show a pronounced increase in interpretability, so much so that observer accuracies far exceed those obtained for late-layer metamers for the standard network.
A second point of stark contrast between the two methods is that of average distances from the reference natural images from which both metamers and SnS invariant images were generated (Table S4). Indeed, metamers are explicitly optimized to minimize representational distance to the target image at a given layer, whereas SnS invariances explicitly maximize that distance. As a result, metamers lie much closer to their reference images than SnS invariances at all depths and for both standard and robust generator networks. For comparison, we also report the average pairwise distances between natural ImageNet images from the same class at the same three layers (i.e., within-category distance for the three representation stages, see Supplementary Section A.5.2). While metamers lie far below these natural-image baselines, SnS invariances systematically exceed them by a wide margin (Table S4). In the context of the different image synthesis between the two methods, we remind that, as detailed in Supplementary Section D, metamer-like stimuli arise as a special case of the SnS objective (i.e. the squeeze-down single loss variant in Fig. S6). In this sense, SnS provides a gradient-free generalization of metamer-like synthesis, enabling analogous invariance characterizations in settings - such as in vivo experiments - where gradients are not accessible.
Table S3: Interpretability comparison between metamers and SnS hierarchical invariances. Classification accuracies (mean $\pm$ SEM) of the standard and $L_{2}$-robust observer networks used in the interpretability experiment (Fig. 4c), evaluated on images generated either as metamers or using SnS with both standard and $L_{2}$-robust ResNet50 as the generator network (Gen Net column). Metrics are reported for the three different representation layers used also in Fig. 4 (Net Layer column). Note that accuracy results for the metamers conditions are in line with Feather et al. (2023).
Condition Gen Net Net Layer Standard observers Robust observers
Metamer Standard Pixel space $0.999\pm 0.00122$ $0.989\pm 0.00653$
Metamer Standard Mid layer $0.939\pm 0.01878$ $0.883\pm 0.05348$
Metamer Standard Late layer $0.310\pm 0.02817$ $0.121\pm 0.01388$
Metamer L2 Pixel space $0.998\pm 0.00122$ $0.989\pm 0.00653$
Metamer L2 Mid layer $0.997\pm 0.00163$ $0.985\pm 0.00816$
Metamer L2 Late layer $94\pm 0.00245 0.990$$90\pm 0.00694$
SnS Standard Pixel space $0.397\pm 0.087$ $0.224\pm 0.014$
SnS Standard Mid layer $0.476\pm 0.099$ $0.256\pm 0.022$
SnS Standard Late layer $0.779\pm 0.070$ $0.636\pm 0.024$
SnS L2 Pixel space $0.843\pm 0.052$ $0.936\pm 0.040$
SnS L2 Mid layer $0.722\pm 0.035$ $0.857\pm 0.041$
SnS L2 Late layer $0.721\pm 0.030$ $0.888\pm 0.009$
Table S4: Comparison of average distances with the reference natural image between metamers and SnS invariances. Mean $\pm$ SEM are reported for each representation layer used in the experiments (Pixel space, Mid layer, and Late layer of ResNet50 standard and $L_{2}$-robust variants). Distances between natural images belonging to the same ImageNet category are reported as reference. For every network layer, SnS invariances yield substantially larger distances than both the natural-image baseline and the metamers.
Condition Net Layer ResNet50 ResNet50 ($L_{2}$)
Metamer Pixel space $0.7998\pm 0.01523$ $0.7998\pm 0.01523$
Metamer Mid layer $978.2\pm 6.483$ $6.181\pm 0.5169$
Metamer Late layer $356.8\pm 10.48$ $8.254\pm 0.9402$
SnS Pixel space $281.4\pm 5.991$ $281.8\pm 4.94$
SnS Mid layer $1935\pm 8.864$ $371.8\pm 2.603$
SnS Late layer $858.2\pm 22.75$ $406.3\pm 6.442$
Natural Pixel space $143.4\pm 1.447$ $142.4\pm 1.669$
Natural Mid layer $1478\pm 8.138$ $201.3\pm 2.268$
Natural Late layer $384.8\pm 7.128$ $239.1\pm 1.845$
Appendix J SnS discovers layer specific hierarchical invariances for $L_{\infty}$-robust models
To assess whether hierarchical divergence in SnS invariances also arises under $L_{\infty}$ robustification, we generated SnS invariances for an $L_{\infty}$-robust ResNet50 ($\epsilon=4/255$) from RobustBench (Croce et al., 2020) using the same n = 120 natural images and the same three stretching depths (pixel-space, mid-level, high-level) as in Fig. 4. Qualitatively (Fig. S11a) and quantitatively (Fig. S11b), layer-specific invariances in the $L_{\infty}$-robust model remained highly distinct and discriminable in pixel space, mirroring those already shown in Fig. 3. We then tested interpretability across observers by presenting the $L_{\infty}$-robust invariances to three architectures (ResNet-18, Wide-ResNet50-2, ConvNeXt-Base), each in standard (Paszke et al., 2019) and $L_{\infty}$-robust versions ($\epsilon=4/255$, also from RobustBench (Croce et al., 2020)). As in the $L_{2}$ setting, invariances from the $L_{\infty}$-robust generator were more interpretable on average than those from a standard ResNet50, and interpretability again decreased when robust invariances were evaluated by standard networks (Fig. S11c). Notably, however, when $L_{\infty}$ invariances were judged by other $L_{\infty}$-robust observers, classification accuracy either remained at ceiling or increased across stretching depths, suggesting that $L_{\infty}$-robust networks share a highly aligned invariance structure even at deeper layers. This observation highlights promising directions for future work comparing invariances across robustness regimes, as well as in models whose robustness arises from perceptually aligned training strategies (e.g., harmonized neural networks (Fel et al., 2022)) or biologically inspired architectures such as VOneResNet (Dapello et al., 2020).
Figure S11: Invariances in $L_{\infty}$-robust ResNet50. (a) Example invariant images obtained for a readout neuron ("cup") in $L_{\infty}$-robust ResNet50 and computed for Pixel space, Layer3_conv1, Layer4_conv7 (rows). (b) Accuracy of a SVC in discriminating the three classes of invariant images produced by stretching at the aforementioned layers as a function of the number of principal components fed to the classifier. (c) Classification accuracies for $L_{\infty}$-robust and standard ResNet50 invariances across multiple $L_{\infty}$-robust and standard observer networks.